### v4

# TIB-STC: A Large-Scale Structured Tibetan Benchmark for Low-Resource Language Modeling

Abstract
Advancement of large language models (LLMs) has brought transformative capabilities to NLP, but such progress remains unevenly distributed, especially for low-resource and culturally rich languages like Tibetan. In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain benchmark specifically designed to support the development and evaluation of LLMs for the Tibetan language. Spanning over 11 billion tokens across literature, religion, medicine, law, and daily communication, TIB-STC preserves traditional grammar and stylistic richness. To validate its utility, we train a reference model, Sun-Shine, on TIB-STC through a three-stage pipeline involving pretraining, supervised fine-tuning, and preference optimization. Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the benchmark’s effectiveness in enabling robust instruction-following and culturally aligned generation. We release TIB-STC to advance research in low-resource language modeling and promote inclusivity in multilingual NLP. All data are available: https://github.com/Vicentvankor/sun-shine.
Introduction
Large language models (LLMs) have demonstrated impressive performance across a broad range of natural language processing (NLP) tasks, exhibiting strong generalization and emergent capabilities (Minaee et al. 2024). However, these advancements have primarily benefited high-resource languages such as English and Chinese (Achiam et al. 2023; Yang and et al. 2024; Touvron et al. 2023), while low-resource and linguistically complex languages—such as Tibetan—remain significantly underserved (Liang et al. 2024).
Tibetan presents unique challenges for NLP, including complex morphological structures, vertically composed syllables, and deep integration with cultural semantics (UCLA International Institute 2014; Khysru et al. 2021). These features, combined with a severe lack of large-scale, domain-diverse, and annotated corpora, have hindered the development of effective LLMs for Tibetan (Gao et al. 2025).
Moreover, mainstream LLMs are typically trained on corpora heavily biased toward dominant languages and cultures (Üstün et al. 2024), limiting their generalization to minority languages. As a result, they often produce outputs that are inaccurate, unsafe, or culturally inappropriate when applied to underrepresented languages (Chao et al. 2024). This underscores the urgent need for high-quality, task-aligned benchmarks that support robust and culturally informed modeling for low-resource languages (Brown et al. 2020; Blodgett et al. 2020).
To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus benchmark. Covering over 11 billion tokens from domains such as literature, law, religion, medicine, and daily communication, TIB-STC is designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization. To validate the benchmark’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline. While the primary contribution of this work lies in the construction of the TIB-STC dataset, Sun-Shine serves as a proof of concept for building culturally aligned and linguistically robust LLMs for Tibetan.
Our main contributions are summarized as follows:
  • We introduce TIB-STC, the first large-scale and structured Tibetan corpus benchmark, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
  • We have constructed four task-specific sub-benchmarks (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
  • We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the benchmark’s effectiveness in instruction-following and safety-aware generation.
  • We evaluate our approach on the TLUE benchmark, a Tibetan adaptation of multilingual evaluation datasets, manually corrected for linguistic and cultural fidelity, highlighting the feasibility of structured benchmarking for minority languages.
Related Work
Tibetan Language Processing
Tibetan remains one of the least-resourced languages in NLP due to its complex grammatical structure, two-dimensional syllabic composition, and limited digitized resources (Liu et al. 2024). The scarcity of Tibetan corpora has severely hindered the development of both traditional NLP tools and modern deep learning models (Liu et al. 2025a, b). Existing work in Tibetan NLP has primarily focused on morphological segmentation, part-of-speech tagging, and rule-based translation, often relying on small, domain-specific datasets or handcrafted features (Liang et al. 2024). However, these approaches lack scalability and generalizability, particularly in the context of large-scale pretraining. Moreover, the diverse functional domains of Tibetan, spanning religious scripture, legal and medical texts, and everyday communication, necessitate a corpus that captures both linguistic richness and cultural depth (Khysru et al. 2021). To date, no large-scale, structured benchmark exists to systematically support foundation model training or evaluation for Tibetan.
Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan benchmark designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
Large Language Models
LLMs such as GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023; Dubey et al. 2024), and DeepSeek (Bi and et al. 2024; Guo and et al. 2025) have demonstrated impressive zero-shot and few-shot performance on a variety of tasks in high-resource languages. However, their effectiveness diminishes considerably in low-resource or culturally divergent settings. Recent efforts have explored domain- or region-specific LLMs by integrating language-specific corpora and instruction-tuning datasets (Taori et al. 2023b; GLM et al. 2024; Wang et al. 2023), but most of these focus on widely spoken or contemporary languages. Research on LLMs for minority languages with non-Latin scripts and typologically distinct grammars, such as Tibetan, remains extremely limited (Üstün et al. 2024; Minaee et al. 2024). In addition, safety alignment and bias mitigation for underrepresented languages are underexplored, despite the increased risk of harmful, inaccurate, or culturally insensitive outputs.
To address these limitations, we present Sun-Shine, a Tibetan LLM trained on TIB-STC, to empirically validate the effectiveness of culturally grounded data in enabling robust and responsible modeling for low-resource language contexts.
TIB-STC
TIB-STC, the first comprehensive and detailed Tibetan corpus benchmark for LLM, consists of five categories of data as shown in the benchmark branch of Table 1. All data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics. For the TIB-STC, Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement. For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
Benchmark Instance Tokens Storage Size
Corpus 122,756,184 11B 18GB
Alpaca-Ti 42,676 76M 109MB
Safety-Prompts-Ti 38,156 62M 116MB
CValues-Ti 2,000 6M 9MB
hh-rlhf-Ti 47,411 98M 154MB
Table 1: Subcategories of TIB-STC and Data Size Statistics
Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances. As shown in Figure 1, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
Figure 1: Two sample texts from the TIB-STC corpus belong to the categories of Buddhism and Culture, respectively.
Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from the Alpaca (Taori et al. 2023b, a) project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table 2.
Domain Category
Daily Conversation Small Talk and Q&A
Education Multidisciplinary Problem and Solving
Coding Code Generation and Debugging
Creative Writing Poetry, Story
Practical Tasks Summary, Translation, Text Proofreading
Scientific Reasoning Logical Reasoning and Calculation
Areas of Expertise Medicine, Law, Finance, etc.
Entertainment Riddles, Jokes, Game Suggestions.
Table 2: Details of Alpaca-Ti
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Figure 2.
Figure 2: A sample from Alpaca-Ti, with “–” denoting a missing input.
Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios and the data size is 38,156. We categorize a total of 15,000 Instruction Attacks into six subtypes, including Goal Hijacking, Prompt Leaking, RolePlay Instructions, Unsafe Instruction Topics, Inquiries with Unsafe Opinions, and Reverse Exposure—each contributing 2,500 instances. In addition, we compile 23,000 samples under Typical Safety Scenarios, covering eight safety-critical topics: Insult (3.6K), Unfairness and Discrimination (3.8K), Crimes and Illegal Activities (3.6K), Physical Harm (1.7K), Mental Health (4K), Privacy and Property (1.3K), Ethics and Morality (1.3K), and Others (4.7K). LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts (Chao et al. 2024; Zheng et al. 2024; Sun et al. 2023). workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Figure 3.
Figure 3: One sample from Safety-Prompts-Ti.
CValues-Ti
CValues-Ti is the benchmark to improve the security of Tibetan language models. Considering work (Moore, Deshpande, and Yang 2024; Rozen et al. 2024; Xu et al. 2023), we propose this branch of a small benchmark because LLM can pose risks or have negative social impacts. For CValues-Ti, it has three categories (safe & responsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work (Xu et al. 2023), we streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Figure 4.
Figure 4: One sample from CValues-Ti.
hh-rlhf-Ti
hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF (Bai et al. 2022; Kirk et al. 2024), but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Figure 5.
Figure 5: One representative sample from the HH-RLHF-Ti dataset, where “…” indicates an interaction between a human and an assistant representing the LLM. During this exchange, the human may modify the content of the question or adjust the phrasing to guide the assistant’s response.
Sun-Shine
Sun-Shine is built on the LLaMA-3 (Touvron et al. 2023; Dubey et al. 2024) architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model (Dubey et al. 2024). Additionally, we integrate NLP model translation with human annotation (Üstün et al. 2024) as supplementary data to enhance Sun-Shine’s accuracy.
Figure 6: The training process of Sun-Shine comprises three key stages: pre-training, supervised fine-tuning (SFT), and direct preference optimization (DPO). Pre-training equips Sun-Shine with a strong foundation in language understanding and generation. SFT further enhances its ability to follow task-specific instructions and produce responses that meet safety and alignment standards. Finally, DPO aligns the model’s outputs with user preferences, improving both relevance and user satisfaction.
As illustrated in Figure 6, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT) (Yu et al. 2021), and Direct Preference Optimization (DPO) (Rafailov et al. 2024). For training details, Table 3 shows the parameters’ settings. Notably, we adopted LoRA-based parameter-efficient fine-tuning (Hu et al. 2021) for both SFT and DPO, which significantly reduced memory usage without sacrificing performance. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
Parameter Training Stage
Pre-Training SFT DPO
Precision bfloat16 bfloat16 bfloat16
Optimizer AdamW AdamW AdamW
Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-6}$
LR Scheduler CosineAnnealing CosineAnnealing CosineAnnealing
Warmup Ratio 0.1 0.1 0.1
Gradient Accumulation 4 4 8
Epochs 3 3 3
Batch Size (tokens) 4096 4096 4096
Max Seq Length 4096 4096 4096
LoRA No Yes (rank=32) Yes (rank=32)
Weight Decay 0.1 0.1 0.1
Dropout 0.1 0.1 0.1
pref_beta 0.1
pref_loss Sigmoid
Table 3: Hyperparameters in Each Stage of Sun-Shine
Pre-Training
During the Pre-training stage, the Sun-Shine is pre-trained on the Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction. And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text. For details,
Supervised Fine-Tuning
Following pretraining, the model undergoes supervised fine-tuning (Yu et al. 2021) with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, to align the model’s behavior with particular tasks or domains while ensuring safety and professionalism.
Direct Preference Optimization
In the stage of DPO (Rafailov et al. 2024; Chao et al. 2024; Zheng et al. 2024), the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
Experimental Result
Evaluation: TLUE Benchmark
We used TLUE (Gao et al. 2025), the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench. Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
Ti-MMLU
Ti-MMLU is based on CMMLU (Li et al. 2024), and the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance. It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark includes 17 STEM tasks, 13 humanities (Human) tasks, 22 social science (Social) tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific (China), meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation framework follows the original CMMLU metrics, including Response Rate (RR), Accuracy (ACC), and Conditional Accuracy (CA). Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
Ti-Safetybench
SafetyBench (Zhang and et al. 2024; Sun et al. 2023) is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness (OFF), Unfairness and Bias (UB), Physical Health (PH), Mental Health (MH), Illegal Activities (IA), Ethics and Morality (EM), and Privacy and Property (PP). Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples. To evaluate safety-related capabilities for Tibetan, TLUE (Gao et al. 2025) employed Ti-Safetybench, a Tibetan adaptation of SafetyBench. The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards. The evaluation methodology incorporates two scoring methods: Direct Answer Calculation (DA) and Concern All Answer Calculation (CAA). DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations. Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
Baseline Model
We compare Sun-Shine with several widely used LLMs, including GPT-4o, GPT-4o-mini, and GPT-3.5-turbo (Achiam et al. 2023), LLaMA-3 18B (Dubey et al. 2024; Touvron et al. 2023), Qwen-2.5 72B (Yang and et al. 2024), and DeepSeek-R1 (Bi and et al. 2024; Guo and et al. 2025).
LLM Temperature Top_p Stream
GPT-3.5-turbo 1.0 1.0 False
GPT-4o 1.0 1.0 False
GPT-o1-mini 1.0 1.0 True
DeepSeek-R1 1.0 None True
LlaMA-3.18B 0.6 0.9 False
Qwen-2.5 0.7 0.8 False
Table 4: Hyperparameters of LLM
(a) Response Rate
(b) Conditional Accuracy
(c) Accuracy
Figure 7: Ti-MMLU: Sun-Shine vs. other LLMs.
Detailed experimental settings and evaluation metrics for each model are provided in Table 4. Specifically, Qwen-2.5 and DeepSeek-R1 were tested via their official APIs, while LLaMA-3 18B was accessed through the LLaMA-API platform. All experimental configurations will be fully documented in the final version to ensure transparency and reproducibility.
Insights from Performance Results
Ti-MMLU
As shown in Figure 7, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs. Among them, Sun-Shine is slightly ahead of these. For the Response Rate, Sun-Shine and GPT-4o achieve the highest Response Rate, with Sun-Shine leading across all sub-branches,shown in Figure 8, demonstrating superior responsiveness to Tibetan prompts enabled by its domain-specific training. For Conditional Accuracy, Sun-Shine and DeepSeek-R1 perform best, with the latter holding a slight advantage. Both models maintain consistently high scores, reflecting strong general knowledge capabilities. In terms of overall Accuracy, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas. This underscores that Sun-Shine’s strong Tibetan understanding stems not only from model design but also from effective use of the TIB-STC dataset and targeted training, allowing an 8B model to surpass the 671B MoE DeepSeek-R1 in key tasks.
(a) The highest total number of indicators.
(b) The second total number of indicators.
Figure 8: Statistics of Sun-Shine ranks first and second based on the number of indicators, across all categories of the Ti-MMLU benchmark.
Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. For more details, please refer to supplementary materials.
Ti-Safetybench
as shown in Figure 9, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
(a) LLMs Across DA
(b) LLMs Across CCA
Figure 9: Ti-SafetyBench: Sun-Shine vs. other LLMs.
Overall, Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
Sun-Shine for Tibetan Culture
As shown in Figure 10, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
Figure 10: Comparison of LLM Classical Chinese Output
(a) Sun-Shine
(b) GPT-4o
(c) DeepSeek-R1
Figure 11: Ancient Poetry Writing: Sun-Shine vs. other LLMs
As shown in Figure 11, we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English. Sun-Shine maintains a neutral tone when addressing common questions, avoiding exaggerated praise or references to Tibetan classics. In contrast, GPT-4o adopts a more amplified and majestic style, reminiscent of divine scripture, while DeepSeek-R1’s responses frequently allude to the bodhisattva Mañjuśrī, appearing slightly ”overfitted” with more refined yet noticeably embellished language than GPT-4o. Overall, Sun-Shine surpasses DeepSeek-R1 and GPT-4o in most performance indicators, with its classical Tibetan style attributed to the curated TIB-STC and training strategy, where human annotation and correction played a key role.
Downstream Task
Translation
we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU (Papineni et al. 2002) and ROUGE (Lin 2004) are used to evaluate this task. As shown in Table 5, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 10 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
LLM BLEU-4 ROUGE-1 ROUGE-2 ROUGE-l
Sun-Shine 42.16 65.78 48.91 51.58
GPT-4o 38.85 61.12 44.33 47.02
DeepSeek-R1 41.85 65.49 47.73 51.04
Qwen-2.5 35.20 58.59 41.25 44.73
Table 5: Comparison of Translation Performance
Tibetan Language Proficiency Test
we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system). As shown in Table 6, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our dataset, TIB-STC, which includes a richer collection of ancient Tibetan texts.
LLM A B C D
Sun-Shine 63 61 44 42
GPT-4o 45 47 53 51
DeepSeek-R1 59 56 67 66
Qwen-2.5 41 43 53 54
Table 6: Comparison of Tibetan Language Proficiency Test
Conclusion
In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC benchmark and the Sun-Shine large language model. Our primary focus is the development of TIB-STC, the first large-scale, high-quality Tibetan corpus designed to address the severe lack of resources in low-resource language research. Covering diverse domains such as literature, religion, news, and dialogue, TIB-STC offers structured, culturally rich data that enables the effective training and evaluation of Tibetan language models. Leveraging this benchmark, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression. Extensive evaluation shows that models trained on TIB-STC outperform existing LLMs across metrics such as accuracy, response rate, and stylistic alignment. These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded benchmarks in advancing low-resource language technologies.
Future Work
As part of our ongoing efforts, we have launched Sun-Shine 2.0, an upgraded version of our Tibetan foundation model with enhanced reasoning capabilities. In parallel, we are curating an extended Tibetan corpus to further improve linguistic and cultural coverage.
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62276055 and 62406062, in part by the Sichuan Science and Technology Program under Grant 2023YFG0288, in part by the Natural Science Foundation of Sichuan Province under Grant 2024NSFSC1476, in part by the National Science and Technology Major Project under Grant 2022ZD0116100, and in part by the Sichuan Provincial Major Science and Technology Project under Grant 2024ZDZX0012.
Appendix A Supplementary Material
Note: S-S = Sun-Shine; G-4o = GPT-4o; L-318B = Llama-3.18B; G-35T = GPT-3.5-Turbo; Q-2.5 = Qwen-2.5; Go1-m = GPT-o1-mini; DK-r1 = DeepSeek-r1.
Details of Response Rate
Category Response Rate
S-S G-4o L-318 G-35T Q-2.5 Go1-m DK-r1
ancient_chinese 0.5366 0.4695 0.1402 0.0732 0.2500 0.2317 0.3232
chinese_civil_service_exam 0.4938 0.4062 0.1688 0.0438 0.2312 0.1000 0.3812
chinese_food_culture 0.5588 0.5147 0.1250 0.0809 0.2279 0.1765 0.4191
chinese_foreign_policy 0.5206 0.4981 0.3458 0.0093 0.2804 0.1589 0.4579
chinese_teacher_qualification 0.5972 0.6089 0.2682 0.0726 0.2458 0.3073 0.4078
elementary_chinese 0.5238 0.4167 0.1667 0.0556 0.2817 0.1667 0.4683
ethnology 0.5481 0.5037 0.1852 0.0963 0.3259 0.2148 0.6148
high_school_politics 0.5105 0.4955 0.1888 0.0979 0.2098 0.1748 0.4685
modern_chinese 0.5431 0.5241 0.1724 0.0862 0.2069 0.2328 0.3707
chinese_literature 0.6127 0.5412 0.1618 0.0833 0.2892 0.2500 0.5000
chinese_history 0.5635 0.4489 0.1672 0.0526 0.2632 0.1734 0.3839
chinese_driving_rule 0.5198 0.4420 0.1985 0.0687 0.2519 0.2595 0.3817
construction_project_management 0.6619 0.4388 0.1942 0.0647 0.2662 0.2230 0.3453
elementary_commonsense 0.5404 0.5354 0.1414 0.0758 0.2828 0.1465 0.4848
traditional_chinese_medicine 0.5459 0.5405 0.2378 0.0703 0.2324 0.2595 0.3027
business_ethics 0.5359 0.5167 0.2057 0.0574 0.2919 0.2249 0.3780
economics 0.5283 0.4465 0.2767 0.0440 0.2516 0.1824 0.3396
education 0.5460 0.5215 0.1840 0.0920 0.2761 0.3129 0.4049
college_education 0.5981 0.5234 0.2336 0.1589 0.4299 0.2617 0.4673
journalism 0.5947 0.5447 0.1744 0.0407 0.3256 0.2209 0.4826
marketing 0.6556 0.5944 0.2278 0.0889 0.2667 0.3500 0.2389
professional_accounting 0.6114 0.5829 0.1829 0.0800 0.1943 0.2400 0.3714
professional_psychology 0.5645 0.5303 0.2198 0.0819 0.1940 0.2586 0.3578
public_relations 0.5920 0.5172 0.1437 0.0862 0.2299 0.2931 0.4598
security_study 0.5704 0.4963 0.1630 0.1185 0.4074 0.3037 0.4000
high_school_geography 0.5678 0.4831 0.2712 0.0678 0.3305 0.1780 0.4237
management 0.5905 0.5857 0.2524 0.0857 0.2429 0.3190 0.3810
sociology 0.5442 0.5177 0.1991 0.0442 0.2832 0.3319 0.3805
marxist_theory 0.6085 0.5556 0.2804 0.0529 0.3122 0.2804 0.4497
college_law 0.4648 0.4519 0.2315 0.0833 0.2870 0.2500 0.3981
global_facts 0.5235 0.4930 0.2282 0.0738 0.3423 0.2349 0.4027
international_law 0.5568 0.5081 0.2270 0.0595 0.3351 0.2541 0.0919
jurisprudence 0.5377 0.5109 0.3187 0.0681 0.3358 0.2409 0.3869
world_religions 0.5000 0.4938 0.1125 0.0938 0.2625 0.1812 0.4688
logical 0.5366 0.4715 0.1463 0.0976 0.2195 0.2195 0.2846
professional_law 0.5071 0.4882 0.2654 0.0474 0.2180 0.2180 0.3033
philosophy 0.5524 0.4667 0.2476 0.0952 0.2190 0.2286 0.0476
world_history 0.5342 0.4907 0.2174 0.0683 0.3354 0.2298 0.4720
arts 0.0750 0.5437 0.1500 0.1000 0.2375 0.2125 0.4500
electrical_engineering 0.5698 0.5349 0.2035 0.1337 0.2965 0.2384 0.3663
actuarial_science 0.3692 0.1887 0.1698 0.1792 0.0642 0.1038 0.5000
engineering_hydrology 0.5755 0.5377 0.2170 0.0660 0.3113 0.3208 0.5189
genetics 0.5455 0.4830 0.1932 0.0795 0.1648 0.1761 0.2216
astronomy 0.5758 0.4606 0.2424 0.0727 0.1455 0.2182 0.3879
college_mathematics 0.5048 0.3238 0.2095 0.1429 0.1619 0.1524 0.1429
college_medical_statistics 0.6038 0.5094 0.1698 0.0566 0.1792 0.1887 0.0660
virology 0.5680 0.5089 0.1479 0.0414 0.1893 0.2899 0.2071
computer_science 0.5686 0.4216 0.1961 0.0882 0.1225 0.1814 0.2696
conceptual_physics 0.5986 0.5918 0.3878 0.0340 0.1565 0.2925 0.2109
anatomy 0.6622 0.5203 0.2027 0.0676 0.1689 0.2635 0.2905
machine_learning 0.5656 0.4754 0.1967 0.0574 0.1639 0.1393 0.2951
high_school_physics 0.6364 0.5455 0.1909 0.0727 0.2455 0.2198 0.3415
high_school_mathematics 0.5915 0.2317 0.1646 0.2012 0.1341 0.1098 0.3415
high_school_chemistry 0.5758 0.3409 0.2955 0.0682 0.1591 0.1288 0.0985
high_school_biology 0.6746 0.4852 0.2604 0.0828 0.1657 0.2249 0.0947
elementary_mathematics 0.5996 0.3870 0.1913 0.1565 0.1696 0.1478 0.5478
legal_and_moral_basis 0.5748 0.5561 0.2523 0.0888 0.3131 0.2523 0.5514
computer_security 0.5789 0.5088 0.2398 0.1053 0.2105 0.2339 0.2164
food_science 0.4895 0.4705 0.1469 0.0909 0.2587 0.1818 0.3077
college_medicine 0.4982 0.4908 0.2674 0.0366 0.1722 0.1538 0.1758
clinical_knowledge 0.5738 0.4557 0.1899 0.1013 0.2447 0.2616 0.1308
professional_medicine 0.5798 0.5505 0.3165 0.0771 0.2527 0.2287 0.2074
human_sexuality 0.5635 0.5317 0.2302 0.0317 0.2778 0.2143 0.3651
agronomy 0.5621 0.5089 0.1775 0.0651 0.2071 0.2485 0.4201
sports_science 0.6424 0.5697 0.2000 0.0909 0.2848 0.2667 0.3879
nutrition 0.5586 0.4759 0.1862 0.0621 0.1172 0.2414 0.2552
elementary_IT 0.5462 0.5378 0.2353 0.1050 0.1975 0.2353 0.2647
Table 7: Response Rate: Sun-Shine vs. Other LLMs
Details of Conditional Accuracy
Category Conditional Accuracy
S-S G-4o L-318 G-35T Q-2.5 Go1-m DK-r1
ancient_chinese 0.2697 0.2500 0.2174 0.2500 0.2195 0.2812 0.2500
chinese_civil_service_exam 0.2562 0.2152 0.2222 0.1429 0.3243 0.2368 0.2658
chinese_food_culture 0.3258 0.3143 0.3529 0.2727 0.3226 0.2143 0.3171
chinese_foreign_policy 0.3444 0.3438 0.2973 0.0000 0.2667 0.3235 0.3448
chinese_teacher_qualification 0.2921 0.2752 0.1667 0.3077 0.3636 0.2571 0.2650
elementary_chinese 0.3126 0.2813 0.2857 0.4286 0.2958 0.2533 0.3145
ethnology 0.3719 0.4109 0.2400 0.2308 0.3864 0.3333 0.5714
high_school_politics 0.3662 0.3225 0.2963 0.2143 0.1667 0.2174 0.3412
modern_chinese 0.3592 0.3265 0.2000 0.4000 0.1250 0.1852 0.3489
chinese_literature 0.2980 0.2000 0.3030 0.1765 0.3220 0.3077 0.2879
chinese_history 0.3407 0.2552 0.2778 0.2941 0.2353 0.2418 0.2543
chinese_driving_rule 0.3299 0.2764 0.1923 0.1111 0.2424 0.3247 0.3721
construction_project_management 0.3470 0.2951 0.2593 0.2222 0.2162 0.2292 0.3253
elementary_commonsense 0.3017 0.2925 0.2500 0.4667 0.1964 0.2759 0.2654
traditional_chinese_medicine 0.2971 0.2100 0.2045 0.2308 0.3023 0.2738 0.2912
business_ethics 0.3981 0.3411 0.2791 0.3333 0.2623 0.2391 0.3802
economics 0.3597 0.3476 0.2955 0.2857 0.1500 0.3341 0.3254
education 0.3667 0.3421 0.3647 0.2667 0.3111 0.1818 0.3845
college_education 0.3125 0.2143 0.3600 0.2353 0.2174 0.2857 0.3478
journalism 0.2981 0.2366 0.2667 0.2857 0.2679 0.2281 0.3098
marketing 0.3768 0.3051 0.2683 0.2500 0.3750 0.3541 0.3254
professional_accounting 0.3748 0.3084 0.3438 0.5714 0.2647 0.3451 0.3746
professional_psychology 0.3985 0.2823 0.2157 0.3158 0.3333 0.3698 0.3945
public_relations 0.2427 0.1889 0.3600 0.1333 0.2250 0.2845 0.4752
security_study 0.2997 0.3333 0.3182 0.3125 0.3636 0.2258 0.3977
high_school_geography 0.4179 0̆.2807 0.2188 0.1250 0.2308 0.2368 0.3627
management 0.3089 0.2913 0.2264 0.3333 0.2941 0.3154 0.5369
sociology 0.3333 0.2997 0.2889 0.1000 0.1562 0.2062 0.3221
marxist_theory 0.3217 0.3143 0.2453 0.1000 0.2712 0.3210 0.3212
college_law 0.3287 0.2842 0.2800 0.2222 0.2258 0.2165 0.3109
global_facts 0.3636 0.3599 0.3235 0.0909 0.2353 0.3000 0.3611
international_law 0.3830 0.2524 0.2381 0.3636 0.2903 0.2714 0.3799
jurisprudence 0.3258 0.2667 0.2519 0.2857 0.2754 0.2786 0.3299
world_religions 0.5975 0.2911 0.2222 0.2667 0.2857 0.2143 0.6475
logical 0.2931 0.2879 0.2778 0.2500 0.2593 0.2849 0.3648
professional_law 0.3178 0.2223 0.1786 0.2000 0.2609 0.2865 0.3802
philosophy 0.3276 0.2429 0.2308 0.3000 0.3478 0.3169 0.3057
world_history 0.5658 0.2907 0.2286 0.5455 0.3148 0.3267 0.4407
arts 0.2759 0.2500 0.3333 0.4375 0.3684 0.2157 0.2453
electrical_engineering 0.3349 0.3478 0.3114 0.1304 0.2745 0.2456 0.2857
engineering_hydrology 0.3684 0.1148 0.1304 0.5714 0.2121 0.2500 0.4730
astronomy 0.3903 0.2237 0.2750 0.3750 0.3333 0.2111 0.3699
medical_statistics 0.2978 0.2656 0.2222 0.5000 0.2105 0.3636 0.3026
computer_science 0.4013 0.2845 0.2500 0.2778 0.3600 0.3966 0.3923
conceptual_physics 0.2729 0.2473 0.2712 0.2000 0.2174 0.2632 0.2799
anatomy 0.2410 0.2338 0.1333 0.2000 0.2000 0.1667 0.2367
machine_learning 0.2319 0.2241 0.2083 0.1429 0.2000 0.2498 0.2729
high_school_physics 0.2667 0.1857 0.1429 0.1250 0.2593 0.2441 0.6366
high_school_mathematics 0.6667 0.5000 0.3333 0.4500 0.5268 0.5246 0.5241
high_school_chemistry 0.2667 0.1974 0.3590 0.1111 0.2857 0.2459 0.3636
high_school_biology 0.3171 0.2544 0.2727 0.5000 0.2500 0.2321 0.3295
elementary_mathematics 0.3146 0.2519 0.2273 0.2778 0.4615 0.2830 0.2941
actuarial_science 0.3400 0.2500 0.4444 0.1579 0.3214 0.2800 0.4737
genetics 0.2941 0.2555 0.3529 0.2143 0.3793 0.3011 0.3582
college_mathematics 0.5882 0.2642 0.2727 0.2000 0.4941 0.1905 0.4896
virology 0.2083 0.1744 0.1200 0.4286 0.2188 0.3094 0.2157
legal_and_moral_basis 0.3821 0.3277 0.2963 0.1579 0.3284 0.2436 0.3684
computer_security 0.3223 0.2848 0.3171 0.1667 0.3167 0.2750 0.3033
food_science 0.3914 0.3824 0.1905 0.2308 0.2703 0.2250 0.4154
college_medicine 0.3134 0.2500 0.2055 0.2000 0.2128 0.2022 0.3326
clinical_knowledge 0.3448 0.3088 0.2222 0.3333 0.3426 0.2143 0.3248
professional_medicine 0.2431 0.2222 0.2605 0.3448 0.2737 0.2569 0.3433
human_sexuality 0.3134 0.3099 0.1726 0.7500 0.3714 0.3174 0.2989
agronomy 0.2791 0.2558 0.2667 0.0909 0.3714 0.2188 0.2626
sports_science 0.4171 0.2666 0.3936 0.2667 0.4465 0.2969 0.3830
nutrition 0.2319 0.2222 0.1481 0.2222 0.2941 0.3833 0.4412
elementary_IT 0.3185 0.3906 0.2500 0.3600 0.2128 0.2874 0.3452
Table 8: Conditional Accuracy: Sun-Shine vs. Other LLMs
Details of Accuracy
Category Accuracy
S-S G-4o L-318 G-35T Q-2.5 Go1-m DK-r1
ancient_chinese 0.1341 0.1220 0.3050 0.0183 0.0549 0.0549 0.1159
chinese_civil_service_exam 0.1062 0.1000 0.0375 0.0062 0.0750 0.0025 0.1562
chinese_food_culture 0.1765 0.1618 0.0441 0.0221 0.0735 0.0294 0.1838
chinese_foreign_policy 0.2056 0.2028 0.1028 0.0000 0.0748 0.0654 0.1869
chinese_teacher_qualification 0.1676 0.1553 0.0447 0.0223 0.0894 0.0503 0.2458
elementary_chinese 0.1866 0.1695 0.0476 0.0238 0.0833 0.0397 0.1865
ethnology 0.2296 0.2456 0.0444 0.0222 0.1259 0.0519 0.3333
high_school_politics 0.1738 0.1834 0.0559 0.0210 0.0350 0.0210 0.2098
modern_chinese 0.1897 0.1654 0.0345 0.0345 0.0259 0.0603 0.1466
chinese_literature 0.1765 0.1882 0.0490 0.0147 0.0931 0.0686 0.1961
chinese_history 0.1920 0.1146 0.0464 0.0155 0.0619 0.0372 0.1238
chinese_driving_rule 0.1676 0.1492 0.0382 0.0076 0.0611 0.0763 0.1985
construction_project_management 0.2230 0.1295 0.0504 0.0144 0.0576 0.0504 0.1727
elementary_commonsense 0.1966 0.1714 0.0354 0.0354 0.0556 0.0354 0.2424
traditional_chinese_medicine 0.1568 0.1323 0.0486 0.0162 0.0703 0.0703 0.0973
business_ethics 0.2057 0.1852 0.0574 0.0191 0.0766 0.0478 0.1914
economics 0.1195 0.1195 0.0818 0.0126 0.0377 0.0252 0.1824
education 0.1865 0.1795 0.0675 0.0245 0.0859 0.0613 0.1779
college_education 0.1867 0.2100 0.0741 0.0374 0.0935 0.1028 0.2056
journalism 0.1829 0.1676 0.0465 0.0116 0.0872 0.0640 0.2035
marketing 0.1833 0.2000 0.0611 0.0222 0.1000 0.1167 0.1667
professional_accounting 0.1886 0.1714 0.0629 0.0457 0.0514 0.0629 0.1523
professional_psychology 0.1897 0.1736 0.0474 0.0259 0.0647 0.0690 0.1681
public_relations 0.1437 0.0977 0.0517 0.0115 0.0517 0.0862 0.1389
security_study 0.1333 0.2222 0.0519 0.0370 0.1481 0.0593 0.1259
high_school_geography 0.1356 0.2333 0.0593 0.0085 0.0763 0.0508 0.2034
management 0.1810 0.1695 0.0571 0.0286 0.0714 0.1163 0.2143
sociology 0.1593 0.1726 0.0575 0.0044 0.0442 0.0619 0.1681
marxist_theory 0.1957 0.1799 0.0688 0.0053 0.0847 0.0741 0.2222
college_law 0.1574 0.1222 0.0648 0.0185 0.0648 0.0833 0.1759
global_facts 0.1822 0.1621 0.0738 0.0067 0.0805 0.0671 0.1409
international_law 0.1405 0.1965 0.0541 0.0216 0.0973 0.0595 0.0270
jurisprudence 0.1405 0.1998 0.0803 0.0195 0.0925 0.0706 0.1752
world_religions 0.1437 0.1188 0.0250 0.0250 0.0750 0.0312 0.2750
logical 0.1546 0.1382 0.0407 0.0244 0.0569 0.0813 0.1220
professional_law 0.1611 0.1009 0.0474 0.0095 0.0569 0.0474 0.0758
philosophy 0.0667 0.1810 0.0571 0.0286 0.0762 0.0571 0.0095
world_history 0.1543 0.1304 0.0497 0.0373 0.1056 0.0807 0.1577
arts 0.1880 0.1500 0.5000 0.0438 0.0875 0.0438 0.1750
electrical_engineering 0.1429 0.1190 0.0581 0.0174 0.0814 0.0523 0.2209
engineering_hydrology 0.1981 0.1799 0.0283 0.3770 0.0660 0.0755 0.1792
astronomy 0.1273 0.1030 0.0667 0.0242 0.0545 0.0970 0.1697
medical_statistics 0.1604 0.1415 0.0377 0.0283 0.0377 0.0472 0.0283
computer_science 0.1618 0.1275 0.0049 0.0245 0.0441 0.0441 0.1127
conceptual_physics 0.1497 0.1361 0.1020 0.0068 0.0340 0.0748 0.1224
anatomy 0.2230 0.1216 0.0270 0.0135 0.0338 0.0405 0.0878
machine_learning 0.1311 0.1000 0.0410 0.0082 0.0328 0.0328 0.1066
high_school_physics 0.1818 0.1964 0.0273 0.0091 0.0636 0.0455 0.2091
high_school_mathematics 0.1182 0.1453 0.0549 0.0549 0.1341 0.0671 0.2378
high_school_chemistry 0.1136 0.0970 0.1061 0.0076 0.0455 0.0682 0.0303
high_school_biology 0.1716 0.1534 0.0710 0.0414 0.0414 0.0651 0.0178
elementary_mathematics 0.1432 0.1217 0.0435 0.0435 0.0783 0.0478 0.1139
actuarial_science 0.0755 0.0472 0.0755 0.0283 0.0849 0.0660 0.2453
genetics 0.1307 0.1403 0.0682 0.0170 0.0625 0.0739 0.0625
college_mathematics 0.1333 0.0952 0.0571 0.0286 0.0952 0.0286 0.0762
virology 0.1183 0.0888 0.0178 0.0178 0.0414 0.1183 0.0769
legal_and_moral_basis 0.2182 0.2022 0.0748 0.0140 0.1028 0.0654 0.3411
computer_security 0.2047 0.1579 0.0076 0.0175 0.0877 0.0877 0.1228
food_science 0.1366 0.1936 0.0280 0.0210 0.0699 0.0140 0.1329
college_medicine 0.1538 0.1245 0.0549 0.0073 0.0366 0.0366 0.0623
clinical_knowledge 0.1772 0.1565 0.0422 0.0338 0.0844 0.0591 0.0549
professional_medicine 0.1410 0.1230 0.0824 0.0266 0.0691 0.0718 0.0691
human_sexuality 0.1746 0.1667 0.0397 0.0238 0.1032 0.0635 0.1667
agronomy 0.1420 0.1302 0.0473 0.0059 0.0769 0.0533 0.2071
sports_science 0.2242 0.2189 0.0727 0.0242 0.1091 0.0788 0.1333
nutrition 0.1241 0.1103 0.0276 0.0138 0.0345 0.0759 0.1034
elementary_IT 0.1092 0.2102 0.0588 0.0378 0.0420 0.0798 0.1387
Table 9: Accuracy: Sun-Shine vs. Other LLMs

### v5

@@
-TIB-STC: A Large-Scale Structured Tibetan Benchmark for Low-Resource Language Modeling
+TIB-STC: A Large-Scale Structured Tibetan Dataset for Low-Resource Language Modeling
 Abstract
 Advancement of large language models (LLMs) has brought transformative capabilities to NLP, but such progress remains unevenly distributed, especially for low-resource and culturally rich languages like Tibetan.
-In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain benchmark specifically designed to support the development and evaluation of LLMs for the Tibetan language.
+In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain dataset specifically designed to support the development and evaluation of LLMs for the Tibetan language.
 Spanning over 11 billion tokens across literature, religion, medicine, law, and daily communication, TIB-STC preserves traditional grammar and stylistic richness.
 To validate its utility, we train a reference model, Sun-Shine, on TIB-STC through a three-stage pipeline involving pretraining, supervised fine-tuning, and preference optimization.
-Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the benchmark’s effectiveness in enabling robust instruction-following and culturally aligned generation.
+Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the TIB-STC’s effectiveness in enabling robust instruction-following and culturally aligned generation.
 We release TIB-STC to advance research in low-resource language modeling and promote inclusivity in multilingual NLP.
@@
 2024; Touvron et al.
-2023), while low-resource and linguistically complex languages—such as Tibetan—remain significantly underserved (Liang et al.
+2023), while low-resource and linguistically complex languages, such as Tibetan, remain significantly underserved (Liang et al.
 2024).
@@
 2024).
-This underscores the urgent need for high-quality, task-aligned benchmarks that support robust and culturally informed modeling for low-resource languages (Brown et al.
+This underscores the urgent need for high-quality, task-aligned datasets that support robust and culturally informed modeling for low-resource languages (Brown et al.
 2020; Blodgett et al.
 2020).
-To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus benchmark.
+To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus dataset.
 Covering over 11 billion tokens from domains such as literature, law, religion, medicine, and daily communication, TIB-STC is designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
-To validate the benchmark’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline.
+To validate the dataset’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline.
 While the primary contribution of this work lies in the construction of the TIB-STC dataset, Sun-Shine serves as a proof of concept for building culturally aligned and linguistically robust LLMs for Tibetan.
 Our main contributions are summarized as follows:
-  • We introduce TIB-STC, the first large-scale and structured Tibetan corpus benchmark, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
-  • We have constructed four task-specific sub-benchmarks (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
-  • We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the benchmark’s effectiveness in instruction-following and safety-aware generation.
+  • We introduce TIB-STC, the first large-scale and structured Tibetan corpus dataset, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
+  • We have constructed four task-specific sub-datasets (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
+  • We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the dataset’s effectiveness in instruction-following and safety-aware generation.
   • We evaluate our approach on the TLUE benchmark, a Tibetan adaptation of multilingual evaluation datasets, manually corrected for linguistic and cultural fidelity, highlighting the feasibility of structured benchmarking for minority languages.
@@
 2021).
-To date, no large-scale, structured benchmark exists to systematically support foundation model training or evaluation for Tibetan.
-Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan benchmark designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
+To date, no large-scale, structured dataset exists to systematically support foundation model training or evaluation for Tibetan.
+Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan dataset designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
 Large Language Models
@@
 TIB-STC
-TIB-STC, the first comprehensive and detailed Tibetan corpus benchmark for LLM, consists of five categories of data as shown in the benchmark branch of Table 1.
+TIB-STC, the first comprehensive and detailed Tibetan corpus dataset for LLM, consists of five categories of data as shown in the dataset branch of Table 1.
 All data selection and cleaning are conducted by experts in the Tibetan language.
-Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
-For the TIB-STC, Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement.
-For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
-Benchmark Instance Tokens Storage Size
+Unlike modern Tibetan, this dataset preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
+For the TIB-STC, Corpus is used for LLM pre-training, while other datasets focus on optimization and enhancement.
+For these four sub-datasets, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
+Dataset Instance Tokens Storage Size
 Corpus 122,756,184 11B 18GB
@@
 We draw inspiration from the Alpaca (Taori et al.
-2023b, a) project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world.
+2023b, a) project at Stanford University and develop the first Tibetan dataset specifically designed for self-supervised instruction fine-tuning in the world.
 It covers the following topics as shown in Table 2.
@@
 2023).
-workload it entails, we opted to create a benchmark to simulate the process in advance.
+workload it entails, we opted to create a dataset to simulate the process in advance.
 One data sample is shown in Figure 3.
@@
 CValues-Ti
-CValues-Ti is the benchmark to improve the security of Tibetan language models.
+CValues-Ti is the dataset to improve the security of Tibetan language models.
 Considering work (Moore, Deshpande, and Yang 2024; Rozen et al.
 2024; Xu et al.
-2023), we propose this branch of a small benchmark because LLM can pose risks or have negative social impacts.
+2023), we propose this branch of a small dataset because LLM can pose risks or have negative social impacts.
 For CValues-Ti, it has three categories (safe & responsibility = rejection & positive advice; safe = rejection; unsafe = risk response).
@@
 Unlike the work (Xu et al.
-2023), we streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments.
+2023), we streamlined and condensed the instructions, and compared to their dataset (20K), our sub-dataset, CValues-Ti (2K), achieved equivalent results in subsequent experiments.
 One data sample is shown in Figure 4.
@@
 hh-rlhf-Ti
-hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention.
+hh-rlhf-Ti is also the first dataset that improves the model effect by introducing human intervention.
 It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF (Bai et al.
 2022; Kirk et al.
-2024), but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives.
+2024), but we reset and summarize the previous dataset based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives.
 This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan.
@@
 Conclusion
-In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC benchmark and the Sun-Shine large language model.
+In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC dataset and the Sun-Shine large language model.
 Our primary focus is the development of TIB-STC, the first large-scale, high-quality Tibetan corpus designed to address the severe lack of resources in low-resource language research.
 Covering diverse domains such as literature, religion, news, and dialogue, TIB-STC offers structured, culturally rich data that enables the effective training and evaluation of Tibetan language models.
-Leveraging this benchmark, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression.
+Leveraging this dataset, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression.
 Extensive evaluation shows that models trained on TIB-STC outperform existing LLMs across metrics such as accuracy, response rate, and stylistic alignment.
-These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded benchmarks in advancing low-resource language technologies.
+These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded datasets in advancing low-resource language technologies.
 Future Work